Evaluate log 7(1/49).
Answer by scott8148(6628)

the log of a number is the exponent that the base must be raised to, to get the number

in this case, the base (7) is raised to the -2 power ... 1/49=1/(7^2) ... 1/(7^2)=7^(-2) ... so the log is -2